OpenStudy (anonymous):

3. Given that P(A) = 0.4 and P(B) = 0.5, and P(A and B) = 0.16, determine P(B|A). (Points : 0.40 0.50 0.18 0.10

hartnn (hartnn):

know the formula for P(A/B)??

OpenStudy (dumbcow):

P(B|A) = P(AB)/P(A)

OpenStudy (anonymous):

i get confused

hartnn (hartnn):

the formula for P(B/A) is P(A and B)/P(A)....now you have both the values,just plug in and find answer

OpenStudy (anonymous):

more help please

OpenStudy (anonymous):

Probability of A AND B = 0.4 x 0.5 =0.2 Then we can use the formula Pr(B|A) = Pr (A and B) / Pr (A) = 0.2 / 0.4 = 0.5 which is option B. woo we did it

hartnn (hartnn):

but P(A and B) is given nabsz.j...no need to calculate

OpenStudy (anonymous):

oh whoops my bad

OpenStudy (anonymous):

is tht answer wrong or right

OpenStudy (anonymous):

use the formula and recalculate :)

hartnn (hartnn):

still not getting?